# E-3DGS: Gaussian Splatting with Exposure and Motion Events

### v1

@@
Abstract
@@
Estimating Neural Radiance Fields (NeRFs) from images captured under optimal conditions has been extensively explored in the vision community. However, robotic applications often face challenges such as motion blur, insufficient illumination, and high computational overhead, which adversely affect downstream tasks like navigation, inspection, and scene visualization. To address these challenges, we propose E-3DGS, a novel event-based approach that partitions events into motion (from camera or object movement) and exposure (from camera exposure), using the former to handle fast-motion scenes and using the latter to reconstruct grayscale images for high-quality training and optimization of event-based 3D Gaussian Splatting (3DGS). We introduce a novel integration of 3DGS with exposure events for high-quality reconstruction of explicit scene representations. Our versatile framework can operate on motion events alone for 3D reconstruction, enhance quality using exposure events, or adopt a hybrid mode that balances quality and effectiveness by optimizing with initial exposure events followed by high-speed motion events. We also introduce EME-3D, a real-world 3D dataset with exposure events, motion events, camera calibration parameters, and sparse point clouds. Our method is faster and delivers better reconstruction quality than event-based NeRF while being more cost-effective than NeRF methods that combine event and RGB data by using a single event sensor. By combining motion and exposure events, E-3DGS sets a new benchmark for event-based 3D reconstruction with robust performance in challenging conditions and lower hardware demands. The source code and dataset will be available at https://github.com/MasterHow/E-3DGS.
I Introduction
@@
Reconstructing 3D scenes and objects from images has been a fundamental research focus in various vision applications, including robotics navigation [1, 2, 3], virtual reality [4, 5], and scene understanding [6, 7, 8]. In robotic perception, accurate 3D reconstructions are vital for tasks like localization and path planning [1, 2, 3], while in virtual reality (VR) and augmented reality (AR), they underpin ego localization, spatial reasoning, and immersive visualizations [4, 5]. Neural Radiance Field (NeRF) [9] has become a powerful tool for 3D reconstruction, but its limitations in training and rendering efficiency hinder real-time applications. 3D Gaussian Splatting (3DGS) [10] advances NeRF by using explicit 3D Gaussians and tile-based rasterization, enhancing both efficiency and synthesis quality.
@@
Despite the advancements in NeRF [9] and 3DGS [10], it is notable that both methods heavily rely on high-quality training images to achieve accurate 3D reconstructions with RGB cameras [11]. Unfortunately, motion blur is a common challenge in real-world robotics systems, particularly in fast-moving scenarios or low-light conditions where longer exposure times are necessary. This blur often leads to mismatched feature points in COLMAP [12], resulting in inaccuracies in pose calibration and point cloud initialization, and sometimes even causing failures in recovering camera poses, thereby hampering the training of 3DGS.
@@
Event cameras [13], which capture log-intensity changes asynchronously with microsecond-level resolution, provide a promising solution to these challenges. Several approaches [11, 14, 15, 16] have combined event cameras with NeRF to enhance 3D reconstruction in difficult environments. However, achieving real-time high-fidelity rendering remains a significant challenge for these implicit methods. Recent efforts [17, 18, 19] have explored integrating event cameras with 3D Gaussian Splatting (3DGS). However, these approaches overlook the source of event generation, often relying on additional RGB sensors for fusion or learning-based methods [20] to convert events into pseudo-grayscale images to achieve high-quality reconstructions, making it challenging to obtain such results using a single event sensor.
@@
To achieve high-quality 3D reconstruction using only a single event sensor, we combine motion and exposure events to balance quality and efficiency in high-speed scenarios. We propose three operating modes for E-3DGS to address varying scene reconstruction needs:
@@
  • High-Quality Reconstruction Mode: In this mode, a programmable variable aperture gradually increases sensor brightness. The time each pixel reaches target brightness is recorded, generating grayscale images to guide 3DGS for precise low-light and high-dynamic-range reconstruction.
@@
  • Fast Reconstruction Mode: Utilizing a High-Definition (HD) resolution event camera to capture motion events, which are used to supervise 3D Gaussian Splatting (3DGS), enabling rapid 3D reconstruction in high-speed and low-light scenes.
@@
Moreover, to validate the effectiveness of the proposed E-3DGS, we built a hardware acquisition system using a programmable variable aperture and a high-resolution event camera to create EME-3D, the first real-world 3D reconstruction dataset which distinguishes between Exposure and Motion Events. As shown in Tab. I, EME-3D contains nine sequences of $1280{\times}720$ high-resolution event streams, camera calibration parameters, sparse point clouds, and exposure events for reconstructing high-quality grayscale images. Experimental results demonstrate that, compared to event-based NeRF and event-to-grayscale learning-based 3DGS methods, E-3DGS achieves faster reconstruction, higher quality, and greater flexibility in handling diverse scene requirements. In Fast Reconstruction Mode, E-3DGS achieves a PSNR gain of $5.68dB$ over EventNeRF, along with a significantly higher rendering speed ($79.37$ FPS vs. $0.03$ FPS). In the High-Quality Reconstruction Mode, E-3DGS delivers a PSNR increase of $10.89$ dB compared to the event-to-grayscale learning-based 3DGS baseline.
TABLE I: Comparison of Event-based 3D Reconstruction Datasets
@@
  • We are the first to incorporate exposure event information into event-based 3D Gaussian Splatting (3DGS), converting sparse events during exposure into dense intensity frames for high-quality event-based 3D reconstruction.
@@
  • We establish and release EME-3D, the first real-world 3D dataset that contains both exposure and motion events, camera parameters, and sparse 3D point clouds.
@@
  • We propose a hybrid approach that captures exposure events at low speeds and motion events at high speeds, striking a good balance between reconstruction quality and efficiency.
II Related Work
II-A Neural 3D Reconstruction
@@
Traditional 3D reconstruction methods, such as point clouds [22], meshes [23], and voxel grids [24, 25], rely on explicit representations but are limited by their fixed topological structures. NeRFs [9] have gained traction for synthesizing high-quality novel views using MLP-based neural networks and differentiable volume rendering, but they suffer from low rendering efficiency and long training times. More recently, 3D Gaussian Splatting [10] has demonstrated a significant advancement, offering faster convergence and superior rendering quality by employing Gaussian splats. However, all of these methods require clear RGB inputs and struggle with the fast motion or low-light conditions commonly encountered in real-world applications such as robotics and autonomous systems.
II-B Event-based 3D Reconstruction
@@
Event cameras [13], which detect asynchronous brightness changes at the pixel level, offer substantial benefits in high dynamic range and temporally precise environments. Recent works have leveraged the complementary strengths of event streams and traditional RGB frames to enhance 3D scene reconstruction [11, 16, 19]. Incorporating event data into NeRF-based models [9], such as Ev-NeRF [15] and EventNeRF [14], has demonstrated the potential for multi-view consistency in 3D reconstruction. However, NeRFs’ computational demands hinder real-time rendering. To address this, event-based 3D Gaussian Splatting (3DGS) [17, 26] offers a more efficient alternative, accelerating reconstruction while preserving high-quality geometry and appearance. Nevertheless, due to limited texture in event streams, previous event-based 3DGS reconstructions often lack fine-grained detail [17, 26]. Therefore, some studies have attempted to overcome this by using learning-based event-to-image methods [18, 20], or event-image deblur methods [27, 28] to enhance the reconstruction accuracy. However, learning-based approaches require additional computational resources and yield limited quality improvements, while event-image modal fusion increases hardware complexity, size, and cost. In earlier work, EvTemMap [29] distinguished an event-based photography method by converting the precise timing of exposure events into dense grayscale frames, thereby eliminating the need for multiple sensors. Instead of introducing learning-based event-to-image or event-image deblur modules, we propose to explore exposure and motion events in 3D explicit reconstruction, which enhances both the quality and speed of reconstruction from pure event data, particularly beneficial in low-light or high-dynamic-range scenarios.
@@
III Methodology
In this section, we present E-3DGS, an event-based method that partitions events into motion and exposure for high-quality training and optimization of event-based 3D Gaussian Splatting (3DGS), as illustrated in Fig. 1. Our approach utilizes the high temporal resolution of motion events and the rich texture information from exposure events, effectively integrating both capabilities within a single event camera. In the preliminary section (Sec. III-A), we introduce the 3DGS framework and the event camera model. Next, we explain how to map temporal information from exposure events into high-quality grayscale images (Sec. III-B) and detail the overall loss function in Sec. III-C. Finally, we describe the process for collecting real datasets in Sec. III-D.
III-A Preliminary: 3D Gaussian Splatting & Event Model
@@
III-B Temporal-to-Intensity Mapping of Exposure Events
@@
$TR(t)=\begin{cases}f(t),&0\leq t\leq t_{\text{end}}\\ 1.&t>t_{\text{end}}\end{cases}$ (6)
Each pixel triggers an event at a specific time $t^{*}(u)$, forming a temporal matrix. The photocurrent $I_{\max}(u)$, which represents the scene’s intensity, can be derived from the event time $t^{*}(u)$ as follows:
$I_{\max}(x,y)=\frac{\exp((V_{\text{ref}}+V_{\text{thd}})\cdot C_{\text{PD}})-1}{h(t^{*}(u))},$ (7)
@@
After mapping the time information to grayscale values, normalization is performed by adjusting $I_{\max}(x,y)$ to the range $[0,1]$, yielding a high-resolution grayscale image with adaptive dynamic range. The entire process is robust to various forms of degradation, including noise, contrast threshold variability, and pixel anomalies. The resulting temporal-to-intensity mapping of exposure events can further introduce additional constraints to optimize the event-based 3DGS, as detailed in Sec. III-C.
III-C Loss Functions
The loss function guiding the training process consists of two primary components: the motion event loss and the exposure event loss. The motion event loss is crucial for ensuring that the predicted brightness variations align with the motion events captured by the event camera. It is defined as:
@@ III-C Loss Functions
where $\Delta\hat{L}(u_{k})$ represents the predicted logarithmic brightness change, calculated as $\Delta\hat{L}_{k}=\hat{L}(x_{k},t_{k})-\hat{L}(x_{k},t_{k-1})$. The predicted brightness values, $\hat{I}(u_{k},t_{k})$ and $\hat{I}(u_{k},t_{k-1})$, are obtained from the 3DGS model through volumetric rendering. This formulation ensures that the predicted brightness changes are consistent with the real-world motion event stream observed by the event camera, allowing for accurate motion capture under high-motion scenes.
@@
$L_{\text{rgb}}=\frac{1}{N}\sum_{k=1}^{N}\left(I_{\text{pred}}(u_{\text{k}})-I_{\text{gt}}(u_{\text{k}})\right)^{2},$ (10)
@@
$L=\lambda\cdot L_{\text{evs},\text{norm}}+(1-\lambda)\cdot L_{\text{rgb}},$ (11)
@@
III-D EME-3D: Exposure and Motion Events Dataset for 3D Reconstruction

### v2

# E-3DGS: 3D Gaussian Splatting with Exposure and Motion Events

@@
E-3DGS: 3D Gaussian Splatting with Exposure and Motion Events
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
@@
Achieving 3D reconstruction from images captured under optimal conditions has been extensively studied in the vision and imaging fields. However, in real-world scenarios, challenges such as motion blur and insufficient illumination often limit the performance of standard frame-based cameras in delivering high-quality images. To address these limitations, we incorporate a transmittance adjustment device at the hardware level, enabling event cameras to capture both motion and exposure events for diverse 3D reconstruction scenarios. Motion events (triggered by camera or object movement) are collected in fast-motion scenarios when the device is inactive, while exposure events (generated through controlled camera exposure) are captured during slower motion to reconstruct grayscale images for high-quality training and optimization of event-based 3D Gaussian Splatting (3DGS). Our framework supports three modes: High-Quality Reconstruction using exposure events, Fast Reconstruction relying on motion events, and Balanced Hybrid optimizing with initial exposure events followed by high-speed motion events. On the EventNeRF dataset, we demonstrate that exposure events significantly improve fine detail reconstruction compared to motion events and outperform frame-based cameras under challenging conditions such as low illumination and overexposure. Furthermore, we introduce EME-3D, a real-world 3D dataset with exposure events, motion events, camera calibration parameters, and sparse point clouds. Our method achieves faster and higher-quality reconstruction than event-based NeRF and is more cost-effective than methods combining event and RGB data. By combining motion and exposure events, E-3DGS sets a new benchmark for event-based 3D reconstruction with robust performance in challenging conditions and lower hardware demands. The source code and dataset will be available at https://github.com/MasterHow/E-3DGS.
1 Introduction
@@
Reconstructing 3D scenes and objects from images has been a fundamental research focus in various vision applications, including imaging systems [1, 2], depth perception [3], virtual reality [4, 5], and high-precision measurements [6, 7]. Neural Radiance Field (NeRF) [8] has become a powerful tool for 3D reconstruction, but its limited training and rendering efficiency hinder real-time applications. 3D Gaussian Splatting (3DGS) [9] advances NeRF by using explicit 3D Gaussians and tile-based rasterization, enhancing both efficiency and synthesis quality.
@@
Despite the advancements in NeRF [8] and 3DGS [9], it is notable that both methods heavily rely on high-quality training images to achieve accurate 3D reconstructions with RGB cameras [10]. Unfortunately, motion blur is a common challenge in real-world systems, particularly in fast-moving scenarios or low-light conditions where longer exposure times are necessary [10, 11, 12]. This blur often leads to mismatched feature points in COLMAP [13], resulting in inaccuracies in pose calibration and point cloud initialization, and sometimes even causing failures in recovering camera poses, thereby hampering the training of 3DGS.
@@
Traditional frame cameras capture images by integrating light over a fixed exposure time, but they suffer from limited dynamic range, often losing details in highlights or shadows [14, 15]. In contrast, event cameras [11], which capture log-intensity changes asynchronously with high temporal resolution, provide a promising solution to these challenges. For example, in the digital preservation of cultural heritage, event cameras can capture subtle surface details of artifacts under low-light conditions, minimizing the risk of damage from excessive illumination [16]. Furthermore, their high temporal resolution facilitates the rapid acquisition of 3D models, crucial for digitizing valuable assets quickly and efficiently in dynamic environments [4]. Their low power consumption also makes them well-suited for industrial applications like long-term environmental monitoring or remote robotics, minimizing environmental impact [10, 12]. Several approaches [10, 12, 17, 18] have combined event cameras with NeRF to enhance 3D reconstruction in difficult environments. However, achieving real-time high-fidelity rendering remains a significant challenge for these implicit methods. Recent efforts [19, 20, 21] have explored integrating event cameras with 3D Gaussian Splatting (3DGS). However, these approaches overlook the source of event generation, often relying on additional RGB sensors for fusion or learning-based methods [22] to convert events into pseudo-grayscale images to achieve high-quality reconstructions, making it challenging to obtain such results using a single event sensor.
@@
To achieve high-quality 3D reconstruction using only a single event sensor, we combine motion and exposure events to balance quality and efficiency in diverse scenarios. We introduce three operating modes for E-3DGS to address varying scene reconstruction needs:
@@
  • High-Quality Reconstruction Mode: In this mode, a programmable variable aperture gradually increases sensor brightness. The time each pixel reaches target brightness is recorded, generating grayscale images to guide 3DGS for high-quality reconstruction.
@@
  • Fast Reconstruction Mode: Utilizing a High-Definition (HD) resolution event camera to capture motion events for supervising 3DGS in high-speed scenes.
@@
On the EventNeRF dataset [12], we simulate exposure events and map them to grayscale images, demonstrating their advantages ($35.73$ dB vs. $17.29$ dB) over motion events combined with E2VID in the 3DGS pipeline. By incorporating a programmable variable aperture into event cameras, we enhance their ability to capture fine spatial details, albeit at the cost of reduced effectiveness in mitigating motion blur in high-speed scenarios. Additionally, this method of collecting exposure events retains a significantly higher dynamic range compared to conventional RGB cameras. Under extreme lighting conditions, the advantages of exposure events become particularly evident. In low-light environments, RGB cameras fail to capture contours and other spatial details, whereas grayscale images derived from exposure events maintain high image quality. Similarly, in cases of overexposure, RGB cameras suffer from pixel saturation, losing crucial details in high-brightness regions due to their limited dynamic range, whereas exposure event-based grayscale images effectively preserve and recover fine details in these areas. These results demonstrate the potential of incorporating a programmable variable aperture at the hardware level, combined with a temporal mapping method, to overcome the limitations of conventional RGB cameras and motion-event-only approaches, enabling robust 3D reconstruction using a single event sensor in challenging optical scenarios.
Moreover, to verify the effectiveness of the proposed E-3DGS, we established a hardware acquisition system using a programmable variable aperture and a high-resolution event camera to create EME-3D, the first real-world 3D reconstruction dataset which distinguishes between Exposure and Motion Events. As shown in Tab. 1, EME-3D contains nine sequences of $1280{\times}720$ high-resolution event streams, camera calibration parameters, sparse point clouds, and exposure events for reconstructing high-quality grayscale images. Experimental results demonstrate that, compared to event-based NeRF and event-to-grayscale learning-based 3DGS methods, E-3DGS achieves faster reconstruction, higher quality, and greater flexibility in handling diverse scene requirements. In Fast Reconstruction Mode, E-3DGS achieves a PSNR gain of $5.68dB$ over EventNeRF, along with a significantly higher rendering speed ($79.37$ FPS vs. $0.03$ FPS). In the High-Quality Reconstruction Mode, E-3DGS delivers a PSNR increase of $10.89dB$ compared to the event-to-grayscale learning-based 3DGS baseline.
Table 1: Comparison of Event-based 3D Reconstruction Datasets.
@@
  • This work is the first to incorporate exposure event information into event-based 3D Gaussian Splatting (3DGS), converting sparse events during exposure into dense intensity frames for high-quality event-based 3D reconstruction.
@@
  • To the best of our knowledge, EME-3D is the first real-world 3D dataset that contains both exposure and motion events, camera parameters, and sparse 3D point clouds, and we make it publicly available.
@@
  • A hybrid approach is proposed, capturing exposure events at low speeds and motion events at high speeds, effectively balancing reconstruction quality and efficiency.
2 Related work
2.1 Neural 3D Reconstruction
@@
Traditional 3D reconstruction methods, such as point clouds [24], meshes [25], and voxel grids [26, 27], rely on explicit representations but are limited by their fixed topological structures. NeRFs [8] have gained traction for synthesizing high-quality novel views using MLP-based neural networks and differentiable volume rendering, but they suffer from low rendering efficiency and long training times. More recently, 3D Gaussian Splatting (3DGS) [9] has demonstrated a significant advancement, offering faster convergence and superior rendering quality by employing Gaussian splats. However, all of these methods require clear RGB inputs and struggle with the fast motion or low-light conditions commonly encountered in real-world applications.
2.2 Event-based 3D Reconstruction
@@
Event cameras [11], which detect asynchronous brightness changes at the pixel level, offer substantial benefits in high dynamic range and in scenarios involving high-speed motion or rapid dynamic changes. Recent works have leveraged the complementary strengths of event streams and traditional RGB frames to enhance 3D scene reconstruction [10, 18, 21]. Incorporating event data into NeRF-based models [8], such as Ev-NeRF [17] and EventNeRF [12], has demonstrated the potential for multi-view consistency in 3D reconstruction. However, NeRFs’ computational demands hinder real-time rendering. To address this, event-based 3D Gaussian Splatting (3DGS) [19, 28] offers a more efficient alternative, accelerating reconstruction while preserving high-quality geometry and appearance. Nevertheless, due to limited texture in event streams, previous event-based 3DGS reconstructions often lack fine-grained detail [19, 28]. Therefore, some studies have attempted to overcome this by using learning-based event-to-image methods [20, 22], or event-image deblurring methods [29, 30] to enhance the reconstruction accuracy. However, learning-based approaches require additional computational resources and yield limited quality improvements, while event-image modal fusion increases hardware complexity, size, and cost. In earlier work, EvTemMap [31] distinguished an event-based photography method by converting the precise timing of exposure events into dense grayscale frames, thereby eliminating the need for multiple sensors. Instead of introducing learning-based event-to-image or event-image deblurring modules, we propose to explore exposure and motion events in 3D explicit reconstruction, which enhances both the quality and speed of reconstruction from pure event data, particularly beneficial in low-light or high-dynamic-range scenarios.
@@
3 Methodology
In this section, we present E-3DGS, an event-based method that partitions events into motion and exposure for training and optimization of 3D Gaussian Splatting (3DGS), as illustrated in Fig. 1. Exposure events are mapped into grayscale images, which are then processed using COLMAP [13] to perform Structure from Motion (SfM), a method that estimates camera poses and initializes sparse 3D point clouds based on 2D image sequences. This step is necessary for real-world datasets, while in the EventNeRF dataset [12], camera poses are provided, and the initialization is performed randomly. Subsequently, the real-world motion events and grayscale images derived from exposure events are used as supervision signals, enabling 3D reconstruction tailored to different scene requirements through three operational modes. Our approach utilizes the high temporal resolution of motion events and the rich texture information from exposure events, effectively integrating both capabilities within a single event camera. In the preliminary section (Sec. 3.1), we introduce the 3DGS framework and the event camera model. Next, we detail the process of mapping temporal information from exposure events into high-quality grayscale images (Sec. 3.2), the method for selecting the event window (Sec. 3.3), and the formulation of the overall loss function (Sec. 3.4).
3.1 Preliminary: 3D Gaussian Splatting & Event Model
@@
3.2 Temporal-to-Intensity Mapping of Exposure Events
@@
$TR(t)=\begin{cases}f(t),&0\leq t\leq t_{\text{end}}\\ 1,&t>t_{\text{end}}\end{cases}$ (6)
where $t_{\text{end}}$ represents the end time of the adjustment period, and $f(t)$ is a predefined function governing the transition of transmittance over time. Each pixel triggers an event at a specific time $t^{*}(u)$, forming a temporal matrix. The intensity-proportional value $I_{max}(u)$, which represents the scene’s intensity, can be derived from the event time $t^{*}(u)$ as follows:
$I_{max}(u)=\frac{C}{h(t^{*}(u))},$ (7)
@@
where $t^{*}(u)$ represents the timestamp at which the initial positive event (IPE) is triggered for the pixel at $u=(x,y)$. Here, $C$ is a constant representing the threshold charge required to trigger an event, incorporating sensor parameters like the photodiode capacitance ($C_{PD}$), reference voltage ($V_{ref}$), and threshold voltage ($V_{thd}$). After mapping the time information $t^{*}(u)$ to intensity-proportional grayscale values $I_{max}(u)$ using Eq. (7), normalization is performed by adjusting $I_{max}(u)$ to the range $[0,1]$, yielding a high-resolution grayscale image with adaptive dynamic range. This normalization step effectively handles the pixel-specific constant $C$. The process is robust to various forms of degradation (including noise, contrast threshold variability, and pixel anomalies) because its degradation model explicitly simulates and incorporates these specific real-world imperfections during training data generation, thereby enabling the final restoration model to learn how to handle and mitigate them, as detailed in prior work [31]. The resulting temporal-to-intensity mapping of exposure events can further introduce additional constraints to optimize the event-based 3DGS, as detailed in Sec. 3.4.
3.3 Event Window Selection
As shown in Fig. 1, we select two viewpoints to compute the motion event loss, leveraging the events occurring between them as supervision signals. To achieve this, we first define the start and end points of the event windows. Specifically, the event stream is divided into intervals represented by $t_{i}=i/N_{\text{windows}}$, where $i\in{1,\ldots,N_{\text{windows}}}$ and $t_{i}$ denotes the window endpoints. Fixed-length windows pose significant challenges, as short intervals can impede the propagation of global illumination effects, while long intervals may degrade the fidelity of local details. To address this, we employ a randomized window sampling strategy [12]. Given a fixed window end $t$, the window start $t_{0}$ is sampled as $t_{0}\sim\mathcal{U}[t-L_{\text{max}},t)$, or equivalently, the window length is drawn as $t-t_{0}\sim\mathcal{U}(0,L_{\text{max}}]$. This approach enables the model to learn both global lighting and fine-grained details, striking a balance between global coherence and local accuracy during reconstruction.
3.4 Loss Functions
The loss function guiding the training process consists of two primary components: the motion event loss and the exposure event loss. The motion event loss is crucial for ensuring that the predicted brightness variations align with the motion events captured by the event camera. Based on the event window selection process detailed in Sec. 3.3, the motion event loss is formulated using the events occurring within the time interval defined by the start timestamp $t_{k_{0}}$ and the end timestamp $t_{k}$ as:
@@ 3.4 Loss Functions
where $\Delta\hat{L}(u_{k})$ represents the predicted logarithmic brightness change, calculated as $\Delta\hat{L}_{k}=\hat{L}(u_{k},t_{k})-\hat{L}(u_{k},t_{k_{0}})$. The predicted brightness values, $\hat{I}(u_{k},t_{k})$ and $\hat{I}(u_{k},t_{k_{0}})$, are obtained from the 3DGS model through volumetric rendering. This formulation ensures that the predicted brightness changes are consistent with the real-world motion event stream observed by the event camera, allowing for accurate motion capture under high-motion scenes.
@@
$L_{\text{img}}=\frac{1}{N}\sum_{k=1}^{N}\left(I_{\text{pred}}(u_{\text{k}})-I_{\text{gt}}(u_{\text{k}})\right)^{2},$ (10)
@@
$L=\lambda\cdot L_{\text{evs},\text{norm}}+(1-\lambda)\cdot L_{\text{img}},$ (11)
@@
4 Experiments
4.1 Evaluation Dataset
Synthetic EventNeRF Dataset. The EventNeRF [12] dataset provides RGB images and corresponding motion event streams for seven sequences, featuring a constant gray background. To maintain a consistent background intensity across each sequence, we first segment the RGB images using EfficientSAM [33], as our focus is on 3D reconstruction of foreground objects. To realistically simulate varying lighting, we scale the segmented foreground intensities by factors of $0.25$ for low-light and $1.5$ for overexposure. The segmented foreground is then normalized into $100$ intensity levels ranging from 00 to $1$, simulating the aperture’s transition from fully closed to fully open. These intensity-modulated images are processed through DVS-Voltmeter [34] to generate exposure event streams. The resulting event streams are converted into grayscale images using the Temporal-to-Intensity Mapping method described in Section 3.2. Finally, the background is restored to its designated intensity value using the segmentation masks, ensuring photometric consistency across the dataset.
EME-3D Real-World Dataset. We utilize an AT-DVS (Adjustable Transmittance DVS), specifically the Metavision Evaluation Kit 4 HD by Prophesee equipped with an aperture shutter placed in front of the lens, to capture real-world sequences. Fig. 2 illustrates the experimental setup used for data acquisition. The setup incorporates an optical motorized rotating stage, specifically the YXRA100 rotational table (Beijing Yixuan Yunkong Technology Co., Ltd.; product details in [35]), offering a rotational precision of $0.005^{\circ}$. The imaging system features a $9{\sim}50mm$ LensConnect BH Series variable focal length lens (Edmund Optics, product number $53086$, model EL6Z0915UCS-MPWIR), providing a horizontal field of view (FoV) of $9.18^{\circ}$ and a vertical FoV of $7.78^{\circ}$ during data acquisition. Before the experiment, the object is securely affixed to the stage using Blu-Tack to prevent any relative movement during rotation. Each object undergoes two recording phases, with the rotating stage reset to the zero position before each phase. In the first phase, the stage performs a full $360$° rotation to capture motion events. During the second phase, the stage rotates in $1.8$° increments, controlled by a motion controller, to capture exposure events at specific positions.